# A Survey of Secure Semantic Communications

### v1

@@ 4.8.4 Knowledge Distillation
Knowledge distillation leverages a pre-trained, more sophisticated teacher model to instruct the training of a streamlined student model. Throughout the training phase, the student model mimics the predictive behavior of the teacher model, effectively decreasing the model’s size and parameter count while striving to retain the teacher model’s accuracy levels. Based on whether the teacher model is updated concurrently with the student model, the learning frameworks can be categorically divided into offline distillation, online distillation, and self-distillation.
@@
Figure 23: Residual error based knowledge distillation [247].
@@
Lesson 8
@@
Model compression reduces the size and enhances the efficiency of DL models by minimizing computational requirements without compromising accuracy significantly. Its importance is crucial for deploying Deep JSCC models on resource-constrained devices, as it helps decrease communication costs, conserve storage, and strengthen SemCom systems against multi-party cooperation attacks. Model compression can be achieved through model pruning, parameter quantization, low-rank decomposition, or knowledge distillation. Specifically, model pruning can be designed to defense against model inversion attacks [227], membership inference attacks [228, 229], and adversarial attacks [230, 147]. Parameter quantization can be divided into quantization aware training and post training quantization [232, 231], and the applications can be extended to task oriented SemCom [233] and large knowledge base [235, 236]. Low-rank decomposition can be divided into the following sub-problems: sparsity-enabled decomposition [237, 238], decomposition of convolutional kernels [239], decomposition of each layer in deep JSCC models [240], and task-oriented SemCom [241]. Knowledge distillation can be devided into offline [242, 243], online [244, 245], and self distillation [246, 247], and it is effective to defend against adversarial attacks [248, 249].
@@ 4.9 Physical-Layer Security
where $M_{S}$ is the semantic information at the transmitter. $C_{S}$ represents the encrypted semantic features, that is, the observation of semantic eavesdroppers. Perfect semantic secret necessitates that the mutual information between $M$ and $C$ is zero. If the condition of perfect semantic secret is met, semantic eavesdroppers can only recover the original semantic information from the encrypted semantic characteristics through random guessing, as it represents the optimal strategy. The various random factors present in wireless communications can be effectively harnessed to establish the comparative advantages of legitimate semantic transmission and ensure the security of SemCom. The related PLS techniques include beamforming, artificial noise, relay cooperation, Intelligent reflecting Surface (IRS), physical-layer key generation (PLKG), and physical-layer authentication (PLA), which are descried in detail as follows.
@@ 4.9.1 Beamforming
Figure 24: Semantic-bit coexisting communication framework [250].
@@ 4.9.2 Artificial Noise
Cognitive radio network applications: Jiang et al. [255] proposes two secrecy energy efficiency maximization schemes for energy-efficient secure downlink communication in OFDM-based cognitive radio networks with an eavesdropper having multiple antennas. The proposed schemes exploit the instantaneous and statistical CSI of the eavesdropper, respectively, to defend against eavesdropping and improve energy efficiency. The problems are transformed into equivalent subtractive problems and then approximate convex problems using the difference of two convex functions approximation method. Two-tier power allocation algorithms are proposed to achieve $\epsilon$-optimal solutions.
@@ 4.9.3 Relay Cooperation
Semantic text transmission: Luo et al. [258] introduce a SemCom scheme for wireless relay channels called, which uses an Autoencoder for encoding and decoding sentences from the semantic dimension. The Autoencoder provides anti-noise performance for the system. Additionally, a novel semantic forward mode is designed for the relay node to forward semantic information at the semantic level, particularly in scenarios where the source and destination nodes do not share common knowledge. Hu et al. [259] propose the semantic relay to address the challenges of implementing resource-demanding DL-based SemCom on mobile devices with limited computing and storage resources. The semantic relay serves as an edge server to provide SemCom services for both semantic users with rich resources and conventional users with limited resources. Two new transmission protocols are proposed for text transmissions via the semantic relay, and an optimization problem is formulated to maximize the weighted sum-rate of all users. An efficient algorithm is proposed to solve this non-convex problem, and numerical results demonstrate the effectiveness of the proposed algorithm and the superior performance of the semantic relay compared to traditional decode-and-forward relays, especially in the small bandwidth regime.
@@
Figure 25: Semantic-forward for cooperative communications [261].
@@ 4.9.4 Intelligent Reflecting Surface
Figure 26: Unified framework for IRS-aided SemCom systems [265].
@@
Defense against eavesdroppers: Wang et al. [267] propose the use of a simultaneous transmitting and reflecting reconfigurable intelligent surface (STAR-RIS) to achieve privacy protection in a SemCom system. The STAR-RIS is used to enhance signal transmission between a base station and a destination user, while converting the signal to interference for an eavesdropper. Simulation results show that the proposed method outperforms other benchmarks in protecting SemCom privacy, as evidenced by a significantly lower task success rate achieved by the eavesdropper.
@@ 4.9.6 Physical-Layer Authentication
Figure 27: Environment semantics enabled PLA [274].
@@
Lesson 9
PLS leverages the randomness of wireless channels to ensure secure transmission. Shannon’s concept of “perfect secret” extends to “perfect semantic secret” in SemCom, ensuring eavesdroppers can only recover original information through random guessing. Various PLS techniques, including beamforming, artificial noise, relay cooperation, IRS, PLKG, and PLA, can be harnessed to establish secure semantic transmission. Beamforming can be leveraged in semantic-bit coexisting environments [250], massive MIMO applications [251], and varying environmental conditions [252, 253]. Artificial noises can be used to optimal beamforming [254], and are extended to cognitive radio network [255], cellular vehicle-to-everything network [256], and visible light communication applications [257]. Relay cooperation is effective for semantic text transmission [258], Cooperative SemCom [260, 261], and Multimodal SemCom [262]. IRS can be combined with other PLS techniques, such as artificial noise [279], PLKG [270, 271, 272], and PLA [280].
@@ 5.1 Dynamic and Intelligent Data Cleaning
The future of data cleaning technique will evolve from static preprocessing towards dynamic and intelligent methods. Gradually, dynamic cleaning methods that integrate knowledge graphs and DL will facilitate a comprehensive upgrade of SemCom systems in intricate scenarios. Future research endeavors will concentrate on multimodal data integration, real-time semantic analysis, and cross-domain knowledge inference, thereby further advancing the application of SemCom technology in diverse fields such as healthcare, autonomous driving, and smart cities.
@@ 5.2 Explainable Robust Learning
The robustness and interpretability of deep JSCC models are interrelated. On one hand, enhancing the robustness of JSCC models helps it better cope with various noises and outliers, thereby maintaining stable performance. On the other hand, improving its interpretability aids in better analyzing the basis for its decisions, which in turn facilitates the identification and resolution of robustness issues within the deep JSCC model. Through interpretability research, it becomes easier to pinpoint errors in semantic understanding and to carry out targeted optimizations and improvements.
@@ 5.3 Multi-Strategy Combined Backdoor Defense
As technology progresses, the forms and techniques of backdoor attacks are continually evolving. Attackers may employ diverse trigger patterns, injection methods, and attack targets to execute these attacks, making them highly varied and intricate. Consequently, a solitary defense strategy often falls short in adequately addressing the multifaceted nature of backdoor attacks. A defense approach that integrates multiple strategies can comprehensively consider security needs from various perspectives and defend against backdoor attacks from multiple angles. For example, data cleaning ensures the authenticity and integrity of training data, mitigating the impact of tampered or injected malicious data on deep JSCC models. Enhancing the robustness of deep JSCC models and applying pruning and compression techniques can bolster their resistance to malicious attacks, reducing redundancy and the risk of being compromised. Additionally, backdoor detection technology can swiftly identify and eliminate backdoors, thereby ensuring the security of deep JSCC models. Therefore, the defense method that combines multiple strategies not only prioritizes the security of individual models or components but also takes into account the interactions and impacts among various components from the perspective of the entire SemCom system, ultimately enhancing the overall security and stability of SemComs.

### v2

@@ 4.8.4 Knowledge Distillation
Knowledge distillation leverages a pre-trained, more sophisticated teacher model to instruct the training of a streamlined student model. Throughout the training phase, the student model mimics the predictive behavior of the teacher model, effectively decreasing the model’s size and parameter count while striving to retain the teacher model’s accuracy levels. Based on whether the teacher model is updated concurrently with the student model, the learning frameworks can be categorically divided into offline distillation, online distillation, and self-distillation. Knowledge distillation-enabled SemCom has been studied for multi-user scenarios and edge devices.
@@
Knowledge distillation for Multi-user SemCom: Liu et al. [55] focus on a multi-user SemCom system characterized by limited training samples and unforeseen interference. To enhance model generalization and reduce its size, they introduce a system leveraging knowledge distillation, where a Transformer-based encoder-decoder serves as the semantic encoder-decoder, and fully connected neural networks function as the channel encoder-decoder. They analyze four types of knowledge transfer and model compression techniques. Key system and model parameters, such as noise and interference levels, the number of interfering users, and the dimensions of the encoder and decoder, are taken into account. Numerical results indicate that knowledge distillation significantly boosts robustness and generalization against unexpected interference and mitigates performance degradation when the model size is compressed. Nguyen et al. [261] study the downlink communication from a base station to users with varying computational capacities. The users utilize different versions of Swin transformer models for source decoding and a streamlined channel decoding architecture. They introduce an innovative training approach that integrates transfer learning and knowledge distillation to enhance the performance of users with limited computing power.
Federated knowledge distillation for SemCom: Xing et al. [239] introduce FedDistillSC, a distributed SemCom system leveraging federated distillation for image classification tasks. Following local training with both public and private datasets, the system refines the local models of each transmitter through weighted aggregation based on receiver-generated classification probabilities. This approach facilitates knowledge sharing without compromising private data, thereby enhancing each transmitter’s semantic extraction capabilities. Notably, our method avoids the need for uploading or downloading model parameters, conserving bandwidth resources and mitigating data security risks associated with parameter leakage. Considering that mismatched knowledge bases disrupt semantic alignment between transmitters and receivers, leading to significant semantic errors, Lu et al. [262] propose a semantic knowledge base synchronization framework leveraging federated knowledge distillation for knowledge base establishment and dynamic updates. The framework employs a mutual distillation mechanism to extract knowledge from diverse local knowledge bases. Additionally, they compress the global knowledge base to boost synchronization efficiency.
Combination of knowledge distillation and LLMs for SemCom: Nam et al. [240] propose the language-oriented SemCom by merging LLMs and generative models with the SemCom paradigm. They introduce three innovative algorithms: 1) Semantic source compression, which compresses text prompts into key headwords preserving syntactic essence and context; 2) Semantic error mitigation, which enhances robustness by substituting headwords with synonyms to counter errors; and 3) Listener-adaptive prompt generation, which creates customized prompts tailored to the listener’s language style using in-context learning.
Knowledge distillation-enabled SemCom for edge devices: Albaseer et al. [241] introduce a SemCom framework tailored for heterogeneous, resource-constrained edge devices and computation-intensive servers. They leverages dynamic knowledge distillation to customize semantic models for each device, balancing computational and communication constraints while ensuring QoS. They formulate an optimization problem and develop an adaptive algorithm that iteratively refines semantic knowledge on edge devices, resulting in tailored models suited to their resource profiles.
Figure 23: Residual error based knowledge distillation [238], where the Assistant (middle) is employed to enhance the distillation of knowledge from the Teacher (left). The assistant aids in optimizing the transfer of knowledge, with the goal of capturing the residual error between the Student’s (right) and Teacher’s feature maps.
Lesson 7
@@
Model compression reduces the size and enhances the efficiency of DL models by minimizing computational requirements without compromising accuracy significantly. Its importance is crucial for deploying Deep JSCC models on resource-constrained devices, as it helps decrease communication costs, conserve storage, and strengthen SemCom systems against multi-party cooperation attacks. Model compression can be achieved through model pruning, parameter quantization, low-rank decomposition, or knowledge distillation. Specifically, model pruning can be designed to defense against model inversion attacks [232], membership inference attacks [242, 243], and adversarial attacks [244, 149]. Model pruning for SemCom has also been studied [233, 245]. Parameter quantization can be divided into quantization aware training and post training quantization [234, 246], and the applications can be extended to task oriented SemCom [247], large knowledge base [249, 250], and lightweight SemCom without end-to-end training [235]. Low-rank decomposition can be divided into the following sub-problems: sparsity-enabled decomposition [236, 251], decomposition of convolutional kernels [252], decomposition of each layer in deep JSCC models [237], and task-oriented SemCom [253]. Knowledge distillation can be devided into offline [254, 255], online [256, 257], and self distillation [258, 238], and it is effective to defend against adversarial attacks [259, 260]. Knowledge distillation has also been studied for multi-user SemCom [55, 261, 239, 262], combination with LLMs and SemCom [240], and SemCom edge devices [241].
@@ 4.9 Physical-Layer Security
where $M_{S}$ is the semantic information at the transmitter. $C_{S}$ represents the encrypted semantic features, that is, the observation of semantic eavesdroppers. Perfect semantic secret necessitates that the mutual information between $M$ and $C$ is zero. If the condition of perfect semantic secret is met, semantic eavesdroppers can only recover the original semantic information from the encrypted semantic characteristics through random guessing, as it represents the optimal strategy. The various random factors present in wireless communications can be effectively harnessed to establish the comparative advantages of legitimate semantic transmission and ensure the security of SemCom. The related PLS techniques include beamforming, artificial noise, relay cooperation, Intelligent reflecting Surface (IRS), physical-layer key generation (PLKG), and physical-layer authentication (PLA), which are descried in detail as follows. The representative PLS schemes are listed in Tab. LABEL:physicallayersecurity.
Table 13: List of Representative Physical-Layer Security (PLS) Schemes
[…]
@@ 4.9.1 Beamforming
Figure 24: Semantic-bit coexisting communication framework [263], where a single-cell downlink MU-MISO system is considered. The base station has multiple transmit antennas, and each user possesses a single antenna. The users are categorized into two groups: bit-users utilizing BitCom, and sem-users utilizing SemCom.
@@
Integrated sensing and SemCom systems: Yang et al. [266] and Dai et al. [267, 282] study the security issues of beamforming technology in integrated sensing and SemCom systems. Yang et al. [266] explore the allocation of secure resources in a downlink system, accommodating multiple authorized users and potential eavesdroppers. To augment security at the semantic level, semantic information is extracted from the original data prior to transmission. Successful recovery of the received information at the user end necessitates the use of a knowledge base shared with the base station, which is pre-stored. The objective is to maximize the cumulative semantic secrecy rate for all users while ensuring a minimum quality of service for each and maintaining overall sensing performance. They propose an iterative algorithm utilizing the alternating optimization technique. Simulation outcomes highlight the superiority of the proposed algorithm in terms of secure SemCom and reliable detection capabilities. Dai et al. [267] focus on scenarios involving multiple eavesdroppers. Initially, we introduce the concept of secure semantic rate to quantify the dependability of user information acquisition. Subsequently, they formulate an optimization problem aimed at maximizing secure semantic efficiency. To address this problem, they decompose it into two subproblems: beamforming optimization and semantic parameter optimization. For the beamforming subproblem, they employ the Dinkelbach algorithm with predetermined semantic parameters to simplify the objective function and propose an iterative alternating algorithm to solve a series of convex subproblems. For the semantic parameter subproblem, they mathematically demonstrate the monotonic decrease of the objective function in relation to the semantic parameter, given a fixed beamforming vector. Dai et al. [282] focus on scenarios with multiple eavesdroppers and accounting for both perfect and imperfect CSI. Similarly, they split this problem into two subproblems: beamforming optimization and semantic parameter optimization.
@@ 4.9.2 Artificial Noise
Cognitive radio network applications: Jiang et al. [269] propose two secrecy energy efficiency maximization schemes for energy-efficient secure downlink communication in OFDM-based cognitive radio networks with an eavesdropper having multiple antennas. The proposed schemes exploit the instantaneous and statistical CSI of the eavesdropper, respectively, to defend against eavesdropping and improve energy efficiency. The problems are transformed into equivalent subtractive problems and then approximate convex problems using the difference of two convex functions approximation method. Two-tier power allocation algorithms are proposed to achieve $\epsilon$-optimal solutions.
@@ 4.9.3 Relay Cooperation
Semantic text transmission: Luo et al. [270] introduce a SemCom scheme for wireless relay channels, which uses an Autoencoder for encoding and decoding sentences from the semantic dimension. The Autoencoder provides anti-noise performance for the system. Additionally, a novel semantic forward mode is designed for the relay node to forward semantic information at the semantic level, particularly in scenarios where the source and destination nodes do not share common knowledge. Hu et al. [285] propose the semantic relay to address the challenges of implementing resource-demanding DL-based SemCom on mobile devices with limited computing and storage resources. The semantic relay serves as an edge server to provide SemCom services for both semantic users with rich resources and conventional users with limited resources. Two new transmission protocols are proposed for text transmissions via the semantic relay, and an optimization problem is formulated to maximize the weighted sum-rate of all users. An efficient algorithm is proposed to solve this non-convex problem, and numerical results demonstrate the effectiveness of the proposed algorithm and the superior performance of the semantic relay compared to traditional decode-and-forward relays, especially in the small bandwidth regime.
@@
Figure 25: Semantic-forward for cooperative communications [286], where SemComs are employed at both the relay and destination to minimize the payload of the relay-destination link. The deep JSCC algorithm for this relaying systems is applied to enable the destination to recover the source information accurately with the help of semantic information provided by the relay.
@@ 4.9.4 Intelligent Reflecting Surface
Figure 26: Unified framework for IRS-aided SemCom systems [274], where the edge device hosts the feature encoder, while the edge server hosts the classifier. For classification, the edge device uses a DNN-based feature encoder. This feature is then precoded using a linear precoder. Subsequently, it is transmitted over the IRS-assisted wireless channel, resulting in a noisy signal. The edge server employs its classifier determine the estimated classification outcome.
@@
Defense against semantic eavesdroppers: Wang et al. [288] propose the use of a simultaneous transmitting and reflecting reconfigurable intelligent surface (STAR-RIS) to achieve privacy protection in a SemCom system. The STAR-RIS is used to enhance signal transmission between a base station and a destination user, while converting the signal to interference for an eavesdropper. Simulation results show that the proposed method outperforms other benchmarks in protecting SemCom privacy, as evidenced by a significantly lower task success rate achieved by the eavesdropper.
Defense against semantic jamming attacks: Sun et al. [276] introduce a novel concept utilizing multi-functional IRSs to facilitate semantic anti-jamming communication and computing within MEC-supported integrated aerial-ground networks. This framework leverages a semantic transceiver’s inherent resilience and data compression abilities, while multi-functional-IRSs customize the wireless environment through signal reflection, refraction, amplification, and energy harvesting, achieving extensive global coverage, reliable connectivity, and high-speed computing. They formulate a semantic computation rate maximization problem, considering imperfect jammer’s CSI, energy partition constraints for computation offloading, semantic similarity requirements, semantic computation rate targets, and multi-functional-IRS self-sustainability. By transforming the imperfect CSI into a worst-case scenario using a discretization method, They propose a fast-converging monotonic optimization algorithm combined with decoupled second-order cone programming to achieve a globally optimal solution with minimal feasibility evaluations.
@@ 4.9.6 Physical-Layer Authentication
Figure 27: Environment semantics enabled PLA [279], where angle-delay features are extracted for higher interpretability and efficient neural network training. Using the trained EsaNet, physical layer spoofing attacks can be promptly detected in time-varying wireless environments. You Only Look Once (YOLO), an advanced single-stage object detection network, is also employed to swiftly capture angle-delay features from received signals, minimizing data processing overhead and authentication latency. Subsequently, a lightweight neural network is designed to classify without a detection threshold, making it suitable for unknown wireless environments.
@@
Lesson 8
PLS leverages the randomness of wireless channels to ensure secure transmission. Shannon’s concept of “perfect secret” extends to “perfect semantic secret” in SemCom, ensuring eavesdroppers can only recover original information through random guessing. Various PLS techniques, including beamforming, artificial noise, relay cooperation, IRS, PLKG, and PLA, can be harnessed to establish secure semantic transmission. Beamforming can be leveraged in semantic-bit coexisting environments [263], massive MIMO applications [264], varying environmental conditions [265, 281], and integrated sensing and SemCom systems [266, 267, 282]. Artificial noises can be used to optimal beamforming [268], and are extended to cognitive radio network [269], cellular vehicle-to-everything network [283], and visible light communication applications [284]. Relay cooperation is effective for semantic text transmission [270], Cooperative SemCom [271, 286], and Multimodal SemCom [272]. IRS can be combined with other PLS techniques, such as artificial noise [296], PLKG [290, 291, 278], and PLA [297].
@@ 5.1 Dynamic and Intelligent Data Cleaning
The future of data cleaning technique will evolve from static preprocessing towards dynamic and intelligent methods. The primary objective is to ensure semantic consistency and contextual adaptability of data within dynamic, multimodal, and semantic-driven environments. Below are the pivotal challenges and their potential solutions.
Research Challenges: (1) Complexity of semantic noise: In the realm of SemCom, noise is not limited to numerical errors caused by sensor inaccuracies, but also encompasses semantic ambiguities, such as vague textual descriptions or erroneous image annotations, and contextual conflicts, like mismatched video actions and audio descriptions. For instance, in autonomous driving, a camera may incorrectly identify a “stop sign” as a “speed limit sign,” resulting in the SemCom system issuing erroneous instructions. (2) Real-time adaptation to dynamic semantics: Semantic environments evolve over time, such as real-time updates of a patient’s condition in medical diagnosis, rendering traditional static cleaning rules inadequate for adapting to dynamic semantic relationships. For example, in remote surgery, sensor data cleaning must identify and correct physiological signal offsets caused by patient position changes in real-time. (3) Heterogeneous fusion of multimodal data: Multimodal data sources, including texts, images, and sensors, require uniform alignment at the semantic level. However, noise characteristics vary significantly across different modalities. Image noise is pixel-based, while text noise is grammatical. A relevant example is in smart cities, where the semantic correlation between traffic camera data (images) and social media texts (such as “congestion on a certain road segment”) necessitates dynamic verification.
Potential Solutions: (1) Multimodal semantic alignment and cleaning technology: Develop cross-modal attention mechanisms to facilitate joint learning of semantic consistency across multimodal data. For instance, in smart healthcare, aligning pathological images with diagnostic reports can help identify and rectify discrepancies in descriptions, such as unreported tumor areas. (2) Lightweight real-time cleaning framework technology: Implement a hierarchical cleaning architecture leveraging edge computing. Lightweight rule filtering, such as outlier detection is executed on terminal devices, while complex semantic analysis is performed in the cloud. For example, in IIoT, sensor endpoints filter noise data in real-time (like sudden temperature spikes), while the cloud analyzes equipment failure patterns using knowledge graphs.
@@ 5.2 Explainable Robust Learning
Explainable robust learning for SemCom emerges as a pivotal research direction for the future. The endeavor to collaboratively optimize robustness and explainability in deep JSCC encounters numerous hurdles, yet promising avenues exist to overcome them.
Research Challenges: (1) Insufficient model robustness: Traditional JSCC models demonstrate robust resilience against physical channel noise, such as additive white Gaussian noise. However, they lack defensive mechanisms against semantic-level adversarial samples, like semantic perturbations or malicious interference. In real-world communication scenarios, the generalization capability of low-frequency semantic samples is limited, causing model performance to plummet when confronted with outliers, such as rare semantic expressions. (2) Trade-off between explainability and performance: The intricate structure of DNNs obscures the decision-making logic, making it challenging to identify specific issues in semantic understanding, such as encoder bottlenecks or decoder preferences. Furthermore, existing explainability tools, including attention visualization and feature attribution, struggle to quantify the dependency of explanation results on channel conditions in communication contexts.
Potential Solutions: (1) Explainability-driven model optimization: Utilize a phased explainable architecture, such as encoding-transmission-decoding decoupling, combined with attention visualization at the semantic encoding layer, to pinpoint bottlenecks in semantic understanding. Incorporate causal graph models or counterfactual analysis to uncover spurious correlations in semantic decisions, such as pseudo-associations between channel noise and semantic labels. Additionally, construct a semantic verification module based on knowledge graphs to validate the rationality of the model’s decision-making logic through external knowledge bases. (2) Joint optimization framework: Develop a multi-objective optimization framework that dynamically adjusts robust training strategies based on explainability analysis results, for instance, by enhancing the protection of attention regions. During deployment, employ an online learning mechanism coupled with lightweight fine-tuning techniques, to adapt in real-time to variations in channel conditions and shifts in semantic distributions.
@@ 5.3 Multi-Strategy Combined Backdoor Defense
With the progression of technology, the methodologies and techniques employed in backdoor attacks are continually undergoing evolution. Attackers can utilize diverse trigger mechanisms, injection methods, and attack objectives to carry out these attacks, rendering them exceptionally varied and intricate. Consequently, a solitary defense strategy frequently falls short in addressing the multifaceted characteristics of backdoor attacks. A defensive methodology that amalgamates multiple strategies can holistically consider security requirements from various viewpoints and defend against backdoor attacks from diverse angles.
Research Challenges: Attackers have the capability to launch backdoor attacks through a multitude of avenues, including semantic deception and malicious code injection. These diversified attack vectors render it challenging for a single defense strategy to fully counteract. Moreover, despite the exceptional performance of deep JSCC models in semantic comprehension and generation, their intricacy and dependencies introduce security vulnerabilities. Attackers might exploit the model’s internal structure or training data for backdoor attacks, or disrupt its normal functionality by tampering with model parameters.
Potential Solutions: Training the model with a diverse array of datasets can bolster its resilience to data, thereby mitigating the impact of tampered or injected malicious data on model performance. Concurrently, data cleaning techniques can proficiently eliminate or flag malicious data, ensuring the authenticity and integrity of the training data. Additionally, pruning and compressing the deep JSCC model can decrease its complexity and redundancy, thereby fortifying its resistance to malicious attacks. Pruning techniques can eliminate parameters that contribute minimally to the training data, while compression techniques can reduce the model’s size and diminish the risk of it being compromised.